### v1

# in Post-Training and Test-Time Scaling of Large Language Models

Abstract
Recent developments in Large Language Models (LLMs) have shifted from pre-training scaling to post-training and test-time scaling. Across these developments, a key unified paradigm has arisen: Learning from Rewards, where reward signals act as the guiding stars to steer LLM behavior. It has underpinned a wide range of prevalent techniques, such as reinforcement learning (in RLHF, DPO, and GRPO), reward-guided decoding, and post-hoc correction. Crucially, this paradigm enables the transition from passive learning from static data to active learning from dynamic feedback. This endows LLMs with aligned preferences and deep reasoning capabilities. In this survey, we present a comprehensive overview of the paradigm of learning from rewards. We categorize and analyze the strategies under this paradigm across training, inference, and post-inference stages. We further discuss the benchmarks for reward models and the primary applications. Finally we highlight the challenges and future directions. We maintain a paper collection at GitHub repository.
[…]
citnum
Sailing AI by the Stars: A Survey of Learning from Rewards
in Post-Training and Test-Time Scaling of Large Language Models
Xiaobao Wu Nanyang Technological University xiaobao.wu@ntu.edu.sg
1 Introduction
Recent years have witnessed the rapid advancement of Large Language Models (LLMs), such as ChatGPT OpenAI (2023), Claude Anthropic (2025), and Llama Meta (2023, 2024). These models are initially empowered by pre-training scaling Kaplan et al. (2020), which trains LLMs on massive corpora through next-token prediction. While this approach enables broad linguistic and knowledge representations, it suffers from several fundamental limitations: misalignment with human values Bai et al. (2022b); Zhang et al. (2023b); Deshpande et al. (2023), difficulty in adapting to various task objectives Lyu et al. (2023); Wang et al. (2023a), and deficiencies in deep reasoning Mirzadeh et al. (2024); Wu et al. (2024h). As a result, these limitations confine pre-trained models to surface-level tasks, falling short of the long-term goal of robust and general AI. To address these limitations, recent efforts have turned to post-training and test-time scaling, which seek to further refine LLMs after pre-training.
Figure 1: Illustration of the scaling phases of LLMs. The learning-from-rewards paradigm plays a pivotal role in the post-training and test-time scaling.
Across the post-training and test-time scaling, a critical unified paradigm has emerged as illustrated in Figure 1: Learning from Rewards, which leverages reward signals to guide model behavior through diverse learning strategies. For post-training scaling, this paradigm has underpinned several key techniques, including preference alignment through Reinforcement Learning from Human Feedback (RLHF, Ouyang et al., 2022) or AI Feedback (RLAIF, Bai et al., 2022b) with scalar rewards and PPO Schulman et al. (2017), and Direct Preference Optimization (DPO, Rafailov et al., 2023) with implicit rewards. For test-time scaling, this paradigm supports eliciting long Chain-of-Thoughts reasoning via GRPO Shao et al. (2024) with rule-based rewards in DeepSeek-R1 DeepSeek-AI et al. (2025), generate-then-rank (Best-of-N) Cobbe et al. (2021); Lightman et al. (2023), reward-guided decoding Deng and Raffel (2023); Khanov et al. (2024), and post-hoc correction Akyurek et al. (2023); Madaan et al. (2023). Through these techniques, this paradigm enables LLMs to learn actively from dynamic feedback, in contrast to learning passively from static data. As such, this endows LLMs with aligned preferences and deep reasoning and planning abilities, leading to more intelligent agents. In consequence, this paradigm has inspired many applications, such as mathematical reasoning Wang et al. (2023b); DeepSeek-AI (2025), code generation Zhu et al. (2024); Zhou et al. (2025a), multimodality Liu et al. (2025h), agents Xia et al. (2025); OpenAI (2025), and embodied AI Zhang et al. (2025d); Zhao et al. (2025a).
Figure 2: A unified framework of learning from rewards. The language model generates outputs conditioned on the inputs; the reward model evaluates the outputs and provides reward signals based on diverse feedback sources and design choices; the learning strategy leverages the rewards to either fine-tune the language model or refine the outputs. This learning-from-rewards paradigm aims to fulfill preference alignment and task-specific goals. The learning strategy can occur at the training, inference, or post-inference stages.
Due to this growing prevalence, we in this paper comprehensively review the learning from rewards for LLMs. We first introduce a taxonomy that categorizes existing works with a unified conceptual framework regarding reward model design and learning strategies. Then we review representative techniques across three main stages: training with rewards, inference with rewards, and post-inference with rewards. We additionally summarize recent reward model benchmarks and finally conclude by outlining key challenges and promising directions for future research.
2 A Taxonomy of Learning from Rewards for LLMs
We first introduce a unified conceptual framework that captures the key components and interactions to understand learning from rewards systemically. Building upon this framework, we then categorize the primary dimensions along which existing methods vary: (i) What is the source of rewards; (i) How to design the reward model; (i) When to learn from rewards; (i) How to learn from rewards.
2.1 A Unified Conceptual Framework
We present a unified conceptual framework for learning from rewards in Figure 2. It abstracts the key components and interactions involved in learning from rewards for language models. In this framework, the language model generates outputs conditioned on the inputs; the reward model then provides rewards to evaluate the output quality; the learning strategy leverages the reward signals to update the language model or adjusts the outputs.
Language Model.
A language model $\mathcal{M}:\mathcal{X}\rightarrow\mathcal{Y}$ generates an output $\hat{y}\in\mathcal{Y}$ given an input $x\in\mathcal{X}$. This formulation covers a wide range of tasks, such as question answering, summarization, and image captioning.
Reward Model.
A reward model evaluates the quality of an output $\hat{y}$ given an input $x$ and produces a reward signal $r$ that reflects desired properties, such as helpfulness, safety, or task-specific correctness. In different contexts, a reward model may be referred to as a verifier and an evaluator. We emphasize that here we adopt a broad definition of the reward model: it can be model-based or model-free. We will discuss these later.
Learning Strategy.
A learning strategy uses reward signals to adjust the behavior of the language model. Here we consider both the training-based (updating model parameters) and training-free strategies (directly refining model outputs).
Figure 3: Reward Model (RM) design dimensions: (a) Model Architecture (Model-based and Model-free); (b) Reward Format (Scalar, Critique, and Implicit); (c) Scoring Pattern (Pointwise and Pairwise); (d) Reward Granularity (Outcome and Process).
2.2 What is the Source of Rewards?
Reward signals originate from two primary sources: human feedback and automated feedback. Each offers trade-offs in terms of reliability, scalability, and cost. We introduce them respectively as follows.
Human Feedback.
Human feedback provides high-quality reward signals grounded in human judgment and intent. It typically collects human annotations through pairwise comparisons between alternative model outputs, e.g., chosen and rejected responses. The collected preference data can be used to train explicit reward models like RLHF Ouyang et al. (2022) or directly fine-tune the language model like DPO Rafailov et al. (2023). While effective, this approach is resource-intensive and may not scale easily across domains or tasks.
Automated Feedback.
To reduce the cost of human annotations and scale up the reward model training, automated feedback has been increasingly explored as an alternative. The automated feedback mainly includes (i) Self-rewarding, where the language model critiques its own outputs Yuan et al. (2024b); Wang et al. (2024e); (i) Trained Models, such as powerful LLMs following the LLM-as-a-Judge design Bai et al. (2022b); Lee et al. (2023a); (i) Predefined Rules, such as accuracy and format rules used in DeepSeek-R1 Shao et al. (2024); DeepSeek-AI et al. (2025). (i) Knowledge, such as structured knowledge base or Wikipedia Peng et al. (2023); Tian et al. (2023). (i) Tools, such as program compilers and interactive systems Le et al. (2022); Liu et al. (2023). The automated feedback enables scalable reward generation but may introduce limitations in interpretability, generality, and alignment quality.
2.3 How to Design the Reward Model?
Designing the reward model is the central foundation of learning from rewards. As shown in Figure 3, we organize the design space into four key dimensions as follows.
  • Base Architecture. As shown in Figure 3(a), this refers to the base architecture of a reward model: (i) Model-based. A dedicated reward model is trained to evaluate outputs. Variants include scalar reward models (outputting numerical scores), generative reward models (producing critiques), and semi-scalar reward models (combining both). Modern rewards models generally adopt the Transformers framework Liu et al. (2024a). (i) Model-free. Reward signals are directly modeled from the source without an explicit trained model, such as DPO Rafailov et al. (2023) and rule-based rewards of GRPO DeepSeek-AI et al. (2025). In order to align with previous literature, we hereafter refer to the reward model as the model-based by default.
  • Reward Format. As shown in Figure 3(b), this describes the specific format of reward signals: (i) Scalar rewards, numerical scores that quantify the quality of model outputs. They are the most commonly used format due to their simplicity and compatibility with learning strategies such as reinforcement learning. Their limitation lies in the sparsity and interpretability. (i) Critique rewards, natural language feedback that evaluates the quality of outputs Saunders et al. (2022); Kwon et al. (2023), such as “The score of this response is 3 out of 5”. They are more expressive and interpretable than scalar rewards, enabling finer-grained guidance, but they may require additional processing to be used in some learning algorithms. (i) Implicit rewardsare signals implicitly embedded in the source without explicit supervision, such as preference data in DPO Rafailov et al. (2023). This format simplifies the implementation but places more burden on the learning strategies to infer appropriate optimization signals.
  • Scoring Pattern. As shown in Figure 3(c), this dimension determines how responses are scored: (i) Pointwise, assigning a score to each response independently; (i) Pairwise, comparing response pairs and selecting the preferred one. The pairwise scoring can be expressed as a scalar score indicating relative preference or a natural language critique such as “Response 1 is better than Response 2”.
  • Reward Granularity. As shown in Figure 3(d), we identify two kinds of reward granularity: (i) Outcome-level, evaluating the holistic quality of outputs. (i) Process-level, assigning rewards to the intermediate steps within the reasoning process of outputs.
2.4 When to Learn from Rewards?
Learning from rewards can occur at different stages of the language model lifecycle:
  • Training with Rewards. At the training stage, reward signals can be transformed into optimization signals by training algorithms to fine-tune the language model, which is the most extensively explored in the literature. It can support post-training alignment with human preference Ouyang et al. (2022); Bai et al. (2022b) and test-time scaling by eliciting the language models’ deep reasoning capabilities through long Chain-of-Thoughts (CoT) DeepSeek-AI et al. (2025).
  • Inference with Rewards. During inference, reward signals can guide the decoding of model outputs without modifying model parameters. This enables test-time scaling by searching in a larger decoding space, such as Best-of-N and tree search Cobbe et al. (2021); Snell et al. (2025).
  • Post-Inference with Rewards. This stage uses rewards to refine model outputs after generation without modifying model parameters. Post-inference with rewards also supports test-time scaling by iteratively refining the outputs Shinn et al. (2023).
2.5 How to Learn from Rewards?
Various learning approaches have been developed to incorporate reward signals to steer model behavior. These approaches are commonly divided into two types: training-based and training-free.
  • Training-based Strategies. Training-based strategies adopt reward signals to optimize the language model parameters. The optimization mainly depends on Reinforcement Learning (RL) and Supervised Fine-Tuning (SFT). Representative examples include Proximal Policy Optimization (PPO, Schulman et al., 2017; Ouyang et al., 2022), Direct Preference Optimization (DPO, Rafailov et al., 2023; Meng et al., 2024), Group Relative Policy Optimization (GRPO, Shao et al., 2024), and Rejection-Sampling Fine-Tuning (RSFT, Nakano et al., 2021; Yuan et al., 2023a; Dong et al., 2023)
  • Training-free Strategies. Training-free strategies leverage reward signals to guide or refine model outputs without updating the language model parameters. They include generate-then-rank, such as Best-of-N Cobbe et al. (2021); Lightman et al. (2023), reward-guided decoding Deng and Raffel (2023); Khanov et al. (2024), and post-inference correction Shinn et al. (2023); Pan et al. (2023a). These methods offer a relatively lightweight mechanism to improve model outputs; some are highly compatible with different model architectures. They are particularly useful when model fine-tuning is infeasible or computationally expensive.
The above methods provide distinct mechanisms to learn the language model from reward signals. We will review them in detail in the following § 3, 4 and 5.
Figure 4: Illustration of Training with Rewards. Based on the reward model design, we mainly focus on scalar rewards, critique rewards, implicit rewards, rule-based rewards, and process rewards. These reward signals are used to fine-tune the language model through reinforcement learning algorithms or supervised fine-tuning.
3 Training with Rewards
In this section, we introduce the methods that incorporate rewards into the training phase of large language models (LLMs). These methods contribute to post-training scaling for preference alignment and test-time scaling by eliciting long Chain-of-Thoughts (CoT) abilities. We begin with a brief review of the primary training algorithms and then categorize existing methods by reward design: (i) Training with Scalar Rewards; (i) Training with Critique Rewards; (i) Training with Implicit Rewards; (i) Training with Rule-based Rewards; (i) Training with Process Rewards; The first three form the core categories, and the latter two, though conceptually overlapped with the former, are separately presented due to their emerging importance.
The primary training algorithms depend on Reinforcement Learning (RL) or Supervised Fine-Tuning (SFT):
  • REINFORCE Sutton et al. (1999). REINFORCE is a fundamental policy gradient algorithm that optimizes expected rewards by estimating gradients through sampled actions and their observed rewards.
  • PPO Schulman et al. (2017). PPO is a widely used reinforcement learning algorithm. It constrains policy updates through clipped objectives to balance learning speed and stability.
  • DPO Rafailov et al. (2023). DPO is a direct preference optimization method that learns from preference data without explicitly modeling rewards.
  • GRPO Shao et al. (2024). GRPO directly uses the average reward of multiple sampled rollouts as the baseline, which avoids the reward and value model of PPO.
  • REINFORCE++ Hu (2025). REINFORCE++ is a variant of the REINFORCE algorithm that incorporates key techniques from PPO without a critic network.
  • RSFT (Rejection-Sampling Fine-Tuning, Nakano et al., 2021; Yuan et al., 2023a). RSFT samples high-reward data offline to construct training datasets for fine-tuning.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=black, rounded corners, align=center, text centered, minimum width=4em, edge+=darkgray, line width=0.5pt, s sep=8pt, line width=0.5pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, minimum width=19em, fill=gray!10, leaf/.style=font=, align=left, draw=none, inner xsep=8pt, leaf2/.style=text width=10em, font=, draw=none, inner xsep=8pt , where level=1text width=6.5em,align=center,font=, where level=2text width=7.5em, align=center, font=, where level=3text width=18em, align=left,font=, where level=4align=left,font=, [ Training with Rewards, ver [ Training with
Scalar Rewards
§ 3.1, fill=training-fill, draw=training-border [ Scalar Rewards
from Human Feedback, fill=training-fill, draw=training-border [ RLHF Ouyang et al. (2022); Safe RLHF Dai et al. (2023);
Fine-Grained RLHF Wu et al. (2023e); Fact-RLHF Sun et al. (2023);
Skywork-Reward Liu et al. (2024a); ImageReward Xu et al. (2023);
RAHF Liang et al. (2024); LiFT Wang et al. (2024f) ,leaf, fill=training-fill, draw=none ] ] [ Scalar Rewards
from Automated Feedback, fill=training-fill, draw=training-border [ RLAIF Bai et al. (2022b); Self-Taught Wang et al. (2024e);
Dutta et al. (2024); VLM-RLAIF Ahn et al. (2024);
RLTF Liu et al. (2023); RLEF Gehring et al. (2024);
StepCoder Dou et al. (2024); RLEF Gehring et al. (2024)
, leaf, fill=training-fill, draw=none ] ] ] [ Training with
Critique Rewards
§ 3.2, fill=training-fill, draw=training-border [ Auto-J Li et al. (2023a); CompassJudger-1 Cao et al. (2024);
Con-J Ye et al. (2024b); GemRM Mahan et al. (2024);
LLaVA-Critic Xiong et al. (2024); DeepSeek-GRM Liu et al. (2025g);
Critic-RM Yu et al. (2024b); MM-RLHF Zhang et al. (2025f) , leaf, fill=training-fill, draw=none, text width= ] ] [ Training with
Implicit Rewards
§ 3.3, fill=training-fill, draw=training-border [ Implicit Rewards
from Human Feedback, fill=training-fill, draw=training-border [ DPO Rafailov et al. (2023); SimPO Meng et al. (2024);
RLHF-V Yu et al. (2023b); UnifiedRM Wang et al. (2025d);
RAFT Dong et al. (2023); ReST Gulcehre et al. (2023);
RSO Liu et al. (2024b); RRHF Yuan et al. (2023b) , leaf, fill=training-fill, draw=none ] ] [ Implicit Rewards
from Automated Feedback, fill=training-fill, draw=training-border [ Self-Rewarding Yuan et al. (2024b);
Meta-Rewarding Wu et al. (2024a);
SCPO Prasad et al. (2024); Zhang et al. (2025c);
PFPO Jiao et al. (2024a); HA-DPO Zhao et al. (2023);
Tian et al. (2023); FLAME Lin et al. (2024a);
TRICE Qiao et al. (2023); CodeLutra Tao et al. (2024);
, leaf, fill=training-fill, draw=none ] ] ] [ Training with
Rule-based Rewards
§ 3.4, fill=training-fill, draw=training-border [ DeepSeek-Math Shao et al. (2024); DeepSeek-R1 DeepSeek-AI et al. (2025);
DAPO Yu et al. (2025b); Open-R1 Face (2025);
Logic-RL Xie et al. (2025a); Visual-RFT Liu et al. (2025h);
CLS-RL Li et al. (2025c); R1-VL Zhang et al. (2025a);
RefAlign Zhao et al. (2025c) , leaf, fill=training-fill, draw=none, text width= ] ] [ Training with
Process Rewards
§ 3.5, fill=training-fill, draw=training-border [ Process Rewards
from Human Feedback, fill=training-fill, draw=training-border [ Uesato et al. (2022); Lightman et al. (2023) , leaf, fill=training-fill, draw=none ] ] [ Process Rewards
from Automated Feedback, fill=training-fill, draw=training-border [ WizardMath Luo et al. (2023); ActPRM Duan et al. (2025);
Math-Shepherd Wang et al. (2023b); PQM Li and Li (2024);
OmegaPRM Luo et al. (2024); HRM Wang et al. (2025b);
PRIME Cui et al. (2025); OREAL Lyu et al. (2025);
GenPRM Zhao et al. (2025b); R-PRM She et al. (2025);
ThinkPRM Khalifa et al. (2025); M-STAR Liu et al. (2024c) , leaf, fill=training-fill, draw=none ] ] ] ]
Figure 5: Overview of Training with Rewards.
3.1 Training with Scalar Rewards
Training the language model with scalar rewards is the most extensively studied strategy in the literature. Most methods derive scalar rewards by training a dedicated reward model, while some extract rewards directly from the source without training a model (i.e., model-free). We introduce the methods based on the reward source human and automated feedback as follows.
Scalar Rewards from Human Feedback.
Human feedback is a key source for constructing scalar rewards. The most prominent example is Reinforcement Learning from Human Feedback (RLHF) Ziegler et al. (2019); Ouyang et al. (2022); Bai et al. (2022a); Glaese et al. (2022). RLHF trains a scalar reward model on human preference data (pairwise comparisons with chosen and rejected responses). The reward models commonly adopt the Transformer architecture with a value head that outputs scalars, and their training objectives follow the Bradley-Terry loss Bradley and Terry (1952), which maximizes the reward differences between preferred and dispreferred outputs. The trained reward model assigns evaluative scalar scores to the model outputs, serving as a proxy for human judgment. For instance, Skywork-Reward Liu et al. (2024a) is a reward model trained on various high-quality human-labeled preference datasets. With the reward model, RLHF fine-tunes the language model through PPO to align it with human preferences, such as harmlessness and helpfulness.
Various variants have been explored based on RLHF. Safe RLHF Dai et al. (2023) emphasizes safety-centric alignment. Fine-Grained RLHF Wu et al. (2023e) moves beyond holistic feedback to dense, segment-level fine-grained feedback regarding relevance, factual accuracy, and completeness.
Training with scalar rewards has also been extended to multimodal tasks. Fact-RLHF Sun et al. (2023) incorporates factuality-aware scalar rewards to reduce hallucinations of multimodal understanding. Others focus on multimodal generation tasks. HPS and its variant Wu et al. (2023d, c) train a binary classifier as the reward model to evaluate generated image quality. Lee et al. (2023b) trains a multimodal reward model based on human feedback and fine-tunes a diffusion model to improve image generation. ImageReward Xu et al. (2023) introduces a general-purpose reward model for text-to-image generation. It is trained on image preferences from human experts and then fine-tunes a diffusion model via a direct tuning algorithm. RAHF Liang et al. (2024) further enriches the reward signals with scalar scores, heatmaps, and keyword omissions to guide sample filtering for fine-tuning. LiFT Wang et al. (2024f) applies a similar idea to the text-to-video task.
Scalar Rewards from Automated Feedback.
While human feedback offers high-quality supervision, it is expensive and difficult to scale. To overcome this challenge, a growing body of work explores automated feedback as a substitute to provide scalar rewards. A prominent example is Reinforcement Learning from AI Feedback (RLAIF) Bai et al. (2022b). RLAIF uses an LLM as a proxy judge to generate preference data following the idea of LLM-as-a-Judge Zheng et al. (2023); Yu et al. (2025a). Like RLHF, RLAIF trains a scalar reward model on them and then uses it to fine-tune the language model. This pipeline substantially reduces reliance on expensive human annotations. Lee et al. (2023a) have shown that RLAIF can achieve comparable performance with RLHF.
Subsequent studies extend this strategy by incorporating diverse forms of automated feedback. Self-Taught Wang et al. (2024e) prompts an LLM to synthesize contrastive synthetic pairs and uses them to fine-tune a reward model. Dutta et al. (2024) collect GPT-3.5’s binary assessments about predefined code quality questions to train a reward model for code generation. VLM-RLAIF Ahn et al. (2024) trains a reward model with self-generated preference data for video-language tasks. It then uses the reward model to fine-tune a video-language model through PPO. Automated feedback can also come from various tools. RLTF Liu et al. (2023) introduces an online reinforcement learning framework for code language models. It derives multi-granularity scalar rewards from real-time unit test execution results, and these rewards are integrated into a unified loss optimized via the REINFORCE algorithm Sutton et al. (1999) to fine-tune the model. Similarly, StepCoder Dou et al. (2024) designs scalar rewards based on compiler feedback, and it directly fine-tunes LLMs through token-level PPO with these rewards, with optimization restricted to executed code tokens. RLEF Gehring et al. (2024) formulates code generation as a multi-turn decision-making process in an interactive environment. It fine-tunes the language model with the program execution results as scalar rewards through turn-level PPO and supports iterative refinement.
3.2 Training with Critique Rewards
Besides scalar rewards, another line of work explores training with critique rewards. This category commonly uses generative reward models to produce natural language critiques on the given response rather than predicting scalar scores. The generated critiques enjoy flexible scoring patterns: they could be pointwise scoring like “The score of this response is 3 out of 5" or pairwise scoring like “Response 1 is better than Response 2". More importantly, the critiques can include explanations for the scoring and refinement suggestions for later improvements, which facilitates reward density and interpretability.
Due to the above advantages, many studies work on generative reward models to produce critique rewards. Auto-J Li et al. (2023a) generates critiques that support pointwise and pairwise evaluation. It adopts GPT-4 to produce evaluation judgments as the training data. CompassJudger-1 Cao et al. (2024) and Con-J Ye et al. (2024b) follow a similar design. SFR-Judges Wang et al. (2024d) fine-tunes an LLM on the response deduction task to improve its judging ability.
Because of the flexibility of generative models, recent studies concentrate on generalist reward models that support multiple tasks and scoring patterns. For instance, LLaVA-Critic Xiong et al. (2024) is a trained reward model for multiple vision-language tasks, supporting both pointwise and pairwise scoring with natural language explanations. DeepSeek-GRM Liu et al. (2025g) is a generalist generative reward model that covers diverse scoring formats and NLP tasks and produces principles for its scoring. We discuss this promising future direction in § 8.
In particular, generative reward models can work beyond critique generation. For example, GenRM Zhang et al. (2024c) considers reward modeling as a token prediction task. Given a prompt like “Is the answer correct?”, it uses the generation probability of answer indicator tokens, like Yes or No, as the scalar rewards. Mahan et al. (2024) follow a similar approach.
Furthermore, several methods adopt a hybrid structure—semi-scalar reward models that produce both scalar and critique rewards. Early attempts incorporate critiques into the training of a scalar reward model, such as CLoud Ankner et al. (2024), MATH-Minos Gao et al. (2024a), and Ye et al. (2024a). Critic-RM Yu et al. (2024b) is a reward model with two heads, one for generating critiques and another for predicting scalars. It is trained on self-generated critiques of model outputs through a multi-task objective. MM-RLHF Zhang et al. (2025f) extends it to the multimodal field. It trains a semi-scalar reward model on the image, video understanding, and safety and then fine-tunes a multimodal language model with the trained reward model.
3.3 Training with Implicit Rewards
Rather than explicit scalars or critiques that evaluate the model outputs, another line of work adopts implicit rewards for training. The reward signals are not provided directly but are implicitly embedded in the structure of the training data, such as curated preference data. We note that some methods may use a scalar reward model to construct the training data, but the reward model does not participate in fine-tuning language models. Instead, the reward signals for fine-tuning are implicitly encoded in the constructed training data, so we categorize them as training with implicit rewards.
Implicit Rewards from Human Feedback.
A pioneering approach using implicit rewards from human feedback is Direct Preference Optimization (DPO) Rafailov et al. (2023). DPO encodes implicit rewards via the log-likelihood difference between preferred and dispreferred responses and steers the generalization of the language model toward preferred ones. DPO proves that its objective is theoretically equivalent to optimizing the Bradley-Terry loss. As such, DPO effectively reduces complicated RLHF Ouyang et al. (2022) into supervised fine-tuning, dramatically simplifying the alignment pipeline.
Several variants have been proposed based on DPO to further simplify the training or expand its applicability. For instance, SimPO Meng et al. (2024) eliminates the need for a reference model by directly optimizing preference pairs using a more straightforward and reference-free objective, further simplifying DPO. KTO Ethayarajh et al. (2024) models the reward signals implicitly by comparing the log-likelihood difference between the model output and a reference baseline. This difference is then transformed through a nonlinear value function to reflect human-like preferences.
Some studies have extended training with implicit rewards in reasoning and multimodal areas. In mathematical reasoning, Step-DPO Lai et al. (2024) and Full-Step-DPO Xu et al. (2025a) build step-level preference data for DPO training to improve their reasoning abilities. In multimodal areas, mDPO Wang et al. (2024a) proposes a multimodal extension of DPO. It introduces conditional and anchored preference objectives to better leverage visual inputs and preserve the likelihood of preferred responses in multimodal tasks. RLHF-V Yu et al. (2023b) fine-tunes a multimodal LLM through DPO with human preferences regarding trustworthiness. UnifiedRM Wang et al. (2025d) trains a unified multimodal reward model on human preferences for several key tasks, including image/video understanding and generation. Later, it builds preference data through the trained reward model for DPO training. Beyond language models, these techniques are applied to diffusion models due to their simplicity, such as Diffusion-DPO Wallace et al. (2024).
Apart from DPO, another line of work follows a Rejection-Sampling Fine-Tuning (RSFT) scheme. They typically select high-quality samples from a large number of candidate data for supervised fine-tuning on the language model. While reward models are usually used to evaluate the candidates, they are not involved in the actual fine-tuning; thus RSFT methods are within the category of training with implicit rewards. RAFT Dong et al. (2023) trains a reward model on the HH-RLHF dataset Bai et al. (2022a) and uses it to select high-reward samples. The model is then directly fine-tuned on these samples. ReST Gulcehre et al. (2023) adopts an iterative framework: it samples responses from the current model, filters high-quality ones with a reward model, and fine-tunes the model on the filtered set. RSO Liu et al. (2024b) trains a reward model on human preferences and then uses it to generate new preference data via statistical rejection sampling from an estimated optimal policy. The language model is fine-tuned on the new preference data. Especially, some approaches directly depend on the ranking loss. RRHF Yuan et al. (2023b) uses a reward model to rank multiple candidate responses. Then it fine-tunes the language model by jointly optimizing a ranking loss and a SFT loss on these ranked responses.
Implicit Rewards from Automated Feedback.
Implicit rewards can originate from diverse automated feedback as well, such as AI feedback, feedback from external knowledge, and feedback from external tools.
  • AI Feedback. AI feedback is a common source of implicit rewards, including self-rewarding and other trained models. Self-Rewarding Yuan et al. (2024b) leverages the language model to evaluate its own outputs with a prompt like Rate this answer from 1 to 5. Then it constructs preference data with high- and low-rated responses to fine-tune the LLM through iterative DPO. Meta-Rewarding Wu et al. (2024a) builds on this idea: the language model judges its outputs and evaluates the quality of its own judgments. This creates a two-level preference structure and enables joint optimization of acting and judging capabilities. Zhang et al. (2025c) extend self-rewarding to the process-level. It constructs step-wise preference data via self-rewarding to improve the language model’s reasoning ability.
    Instead of direct self-assessment, some methods depend on self-consistency to model implicit rewards. SCPO Prasad et al. (2024) samples multiple responses for each input and selects the consistent responses as preferred ones. Similarly, it then fine-tunes the language model through DPO on these constructed preference data. Similarly, PFPO Jiao et al. (2024a) constructs preference data by verifying model-generated solutions against the test cases derived by self-consistency.
    A list of extended studies focuses on the multimodal field. HA-DPO Zhao et al. (2023) evaluates model responses of image understanding with another trained model as a hallucination detector to produce preference data. With these preference data, it fine-tunes the multimodal language models via DPO to reduce hallucination. Many studies follow a similar way to this, such as POVID Zhou et al. (2024c), AMP Zhang et al. (2024d) and RLAIF-V Yu et al. (2024a), and BPO Pi et al. (2024).
  • Feedback from Knowledge and Tools. External knowledge and tools can provide feedback to model implicit rewards. Tian et al. (2023) construct preference pairs by checking whether model outputs are supported by Wikipedia. These preference data are used to fine-tune a language model toward more factually accurate generations via DPO. FLAME Lin et al. (2024a) follows a similar way. It fine-tunes the language model on factuality-aware preference data via both SFT and DPO to maintain the instruction-following ability. TRICE Qiao et al. (2023) uses tool execution feedback to rank sampled outputs, and the language model is trained with a ranking loss combined with SFT on ranked samples toward accurate and appropriate tool usage. CodeLutra Tao et al. (2024) categorizes generated code into successful and failed samples based on code execution outcomes. It constructs preference data with these samples and fine-tunes the model by combining DPO and SFT objectives. Xiong et al. (2025) leverages SymPy as the verifier to construct mathematical reasoning trajectory pairs for DPO training.
Note that some methods use hybrid automated feedback sources. For instance, RefineCoder Zhou et al. (2025a) incorporates critiques from self-rewarding, Elo ranking, and code execution feedback to build the training data for fine-tuning the language model.
3.4 Training with Rule-based Rewards
Recently, training with rule-based rewards  [^1] While rule-based rewards may overlap with previous scalar or critique rewards, here we present them separately due to their emerging significance. has gained prominence, driven by the success of DeepSeek-R1 DeepSeek-AI et al. (2025). Rule-based rewards are derived by verifying outputs against specific rules, such as format constraints and evaluation metrics. By leveraging reinforcement learning with rule-based rewards, DeepSeek-R1 demonstrates that language models can acquire long Chain-of-Thoughts (long CoT) abilities for test-time scaling. This enables it to solve complex tasks such as mathematics and coding and show self-reflection and self-correction behaviors. This phenomenon, characterized by a sudden emergence of advanced reasoning capabilities, is referred to as RL scaling or the “Aha moment". Note that the literature sometimes refers to rule-based rewards as verifiable rewards/outcomes due to their clean evaluation criteria.
In detail, DeepSeek-R1 DeepSeek-AI et al. (2025) defines two types of rule-based rewards: (i) Accuracy Rewards, whether the output is factually or functionally correct, e.g., the correctness of math solutions or code passing unit tests; (ii) Format Rewards, which ensure the output follows specific structural constraints, such as containing special tokens <think>, </think>, <answer>, and </answer> to encourage long CoT reasoning. With these rule-based rewards, it fine-tunes the language model through the RL algorithm GRPO Shao et al. (2024). GRPO eliminates the dependence on the reward and value model in PPO and the preference data construction in DPO.
Later, many following studies have been proposed. DAPO Yu et al. (2025b) and Open-R1 Face (2025) introduce open-source training frameworks, and some extended GRPO algorithms are introduced Xu et al. (2025b); Zuo et al. (2025); Feng et al. (2025c); Zhang et al. (2025b). Logic-RL Xie et al. (2025a) combines accuracy and format rewards to improve logical reasoning through REINFORCE++ Hu (2025). Visual-RFT Liu et al. (2025h) introduces a set of rule-based rewards for visual tasks, such as Intersection over Union (IoU), confidence-based scoring, and format compliance to fine-tune a vision language model via GRPO. These visual tasks include image classification, reasoning grounding, and object detection. Similarly, CLS-RL Li et al. (2025c) designs rule-based rewards for image classification. R1-VL Zhang et al. (2025a) proposes StepGRPO that extends GRPO into step-level multimodal reasoning.
3.5 Training with Process Rewards
The aforementioned strategies mostly depend on outcome rewards—evaluating the holistic quality of outputs. An emerging line of work focuses on training with process rewards [^2] Similarly, we discuss process rewards separately due to their increasing prevalence although they may overlap with previous scalar or critique rewards. which evaluate the intermediate steps in a model’s reasoning trajectory, such as the steps in mathematical problem solving. As shown in Figure 3(d), these methods commonly employ a Process Reward Model (PRM) to assess the intermediate steps. By providing step-level feedback, these methods enable more fine-grained supervision, which especially benefits complex reasoning tasks where intermediate reasoning directly influences the final result, such as mathematics and code.
Process Rewards from Human Feedback.
Early studies leverage human annotations to train PRMs. For instance, Uesato et al. (2022); Lightman et al. (2023) train PRMs using human annotations on intermediate mathematical reasoning steps. Uesato et al. (2022) then use the trained PRM to fine-tune the language model via reinforcement learning to improve its math reasoning.
Process Rewards from Automated Feedback.
A key limitation of PRMs is their reliance on labor-intensive step-level human annotations. To address this limitation, recent efforts propose incorporating automated feedback to supervise PRMs training at scale.
One major direction leverages strong LLMs to generate step-level annotations. For example, WizardMath Luo et al. (2023) uses GPT-4 to label intermediate steps in math solutions and fine-tunes the language model through step-by-step PPO. ActPRM Duan et al. (2025) similarly uses a strong LLM to annotate step-level correctness of reasoning trajectories, which enables step-by-step DPO fine-tuning.
Alternatively, other methods estimate process rewards without explicit annotations. Math-Shepherd Wang et al. (2023b) applies Monte Carlo estimation to infer step-level scores of mathematical reasoning based on the likelihood of reaching a correct final answer. It uses these estimated annotations to train a PRM and then fine-tunes the language model through step-by-step PPO. Jiao et al. (2024b) follow the same way but use DPO training with sampled reasoning trajectories as preference data. PQM Li and Li (2024) reformulates the PRM training as a Q-value ranking problem. It trains a PRM through a margin-based ranking loss over step pairs, encouraging higher Q-values for steps leading to correct answers. OmegaPRM Luo et al. (2024) adopts an efficient and fully automated method to collect step-level supervision with a novel MCTS algorithm. This significantly reduces the annotation cost and enables to scale up the training of high-quality process reward models. HRM Wang et al. (2025b) also leverages reasoning trajectories from MCTS for training. It evaluates both individual steps and consecutive step pairs to capture multi-step coherence and error recovery. Thus this avoids penalizing an early error regardless of subsequent correction.
Some methods attempt to derive process rewards from outcome rewards. Yuan et al. (2024a) derive process rewards from an outcome reward model as the cumulative token-level log-ratio differences. Building on this idea, PRIME Cui et al. (2025) trains a reward model online using only outcome-level supervision (e.g., answer correctness) on sampled solution rollouts for mathematical problems. The reward model estimates token-level process rewards without requiring step-level annotations. It then fine-tunes the language model with these rewards through PPO (or REINFORCE variants). OREAL Lyu et al. (2025) samples reasoning trajectories and their binary outcome rewards from a fine-tuned policy model. Afterward, it trains a PRM to assign importance scores to each token in the trajectories, where the weighted sum of the scores needs to approximate the outcome rewards.
Figure 6: Illustrations of strategies for Inference with Rewards. (a,b): Generate-then-rank with outcome and process rewards. (c): Reward-guided decoding at the token and step level with search algorithms.
Due to the popularity of generative reward models, generative PRMs have also been introduced. For instance, GenPRM Zhao et al. (2025b) predicts the correctness of a reasoning step by generating CoT reasoning and code verification. GenPRM is trained via SFT on synthesized reasoning trajectories from the MATH dataset, where external LLMs generate the rationales and correctness labels. R-PRM She et al. (2025) and ThinkPRM Khalifa et al. (2025) adopt a comparable generative PRM design by incorporating the reasoning process.
In the multimodal field, M-STAR Liu et al. (2024c) trains a multimodal PRM on estimated step-level scores via Monte Carlo rollouts and then uses the reward model to sample high-quality responses to supervise the subsequent fine-tuning iterations.
4 Inference with Rewards
After the training stage, inference with rewards offers a flexible and lightweight mechanism to adapt and steer the model behavior without modifying model parameters. We identify two primary inference-with-rewards strategies: (i) Generate-then-Rank and (ii) Reward-Guided Decoding. These strategies play a critical role for achieving test-time scaling: They allow the language model to search, reflect, and revise its outputs on the fly.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=black, rounded corners, align=center, text centered, minimum width=4em, edge+=darkgray, line width=0.5pt, s sep=8pt, line width=0.5pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, minimum width=19em, fill=gray!10, leaf/.style=font=, align=left, draw=none, inner xsep=8pt, leaf2/.style=text width=10em, font=, draw=none, inner xsep=8pt , where level=1text width=7em,align=center,font=, where level=2text width=8.5em, align=center, font=, where level=3text width=17.5em, align=left,font=, where level=4align=left,font=, [ Inference with Rewards, ver [ Generate-then-Rank
§ 4.1, fill=inference-fill, draw=inference-border [ Ranking by Outcome Rewards, fill=inference-fill, draw=inference-border [ Cobbe et al. (2021); Uesato et al. (2022);
LEVER Ni et al. (2023); V-STaR Hosseini et al. (2024);
GenRM Zhang et al. (2024c); Fast Best-of-N Sun et al. (2024a) ,leaf, fill=inference-fill, draw=none ] ] [ Ranking by Process Rewards, fill=inference-fill, draw=inference-border [ Lightman et al. (2023); DIVERSE Li et al. (2023b);
Math-Shepherd Wang et al. (2023b); ViLPRM Tu et al. (2025)
VisualPRM Wang et al. (2025c);
CoRe Zhu et al. (2022) , leaf, fill=inference-fill, draw=none ] ] ] [ Reward-Guided Decoding
§ 4.2, fill=inference-fill, draw=inference-border [ Token-level Guidance, fill=inference-fill, draw=inference-border [ RAD Deng and Raffel (2023); ARGS Khanov et al. (2024);
PG-TD Zhang et al. (2023c);
ARM Troshin et al. (2024); FaRMA Rashid et al. (2025) , leaf, fill=inference-fill, draw=none ] ] [ Step-level Guidance, fill=inference-fill, draw=inference-border [ CARDS Li et al. (2024a); GRACE Khalifa et al. (2023);
Xie et al. (2023); Snell et al. (2025); ORPS Yu et al. (2024c);
RSD Liao et al. (2025); Tree-of-Thoughts Yao et al. (2023);
OVM Yu et al. (2023a); RAP Hao et al. (2023);
STILL-1 Jiang et al. (2024); rStar Qi et al. (2024);
ReST-MCTS* Zhang et al. (2024a); LE-MCTS Park et al. (2024);
rStar-Math Guan et al. (2025) , leaf, fill=inference-fill, draw=none ] ] ] ]
Figure 7: Overview of Inference with Rewards.
4.1 Generate-then-Rank
The generate-then-rank approach, usually referred to as Best-of-N, easily scales test-time compute to improve model outputs. It samples a number of candidate responses from the language model, scores them with a reward model, and selects the best one as the final output by ranking or voting Wang et al. (2022). Based on the reward granularity, we distinguish two kinds of methods: (i) ranking by outcome rewards and (ii) ranking by process rewards as shown in Figure 6(a,b).
Ranking by Outcome Rewards.
As shown in Figure 6(a), this method adopts an outcome reward model (ORM) to assess the holistic quality of candidate responses. Early work by Cobbe et al. (2021) trains a binary ORM to evaluate the correctness of candidate math solutions and selects the top-ranked one as the final output. Uesato et al. (2022) adopt the same idea and conduct comprehensive experiments on ranking outputs by ORMs. LEVER Ni et al. (2023) trains a binary classifier as the ORM with code execution results as supervision. During inference, it ranks generated candidate code jointly based on the ORM’s score and the generation probability. V-STaR Hosseini et al. (2024) trains a verifier as the ORM on preference pairs through DPO to rank candidate math/code solutions during inference. Its ORM supports iterative training on dynamically collected preference data, which can progressively improve performance. GenRM Zhang et al. (2024c) follows a generative way. It scores multiple candidate solutions by computing the generation probability of the Yes/No token as rewards and then selects the best solution.
While simple and effective, this strategy becomes computationally expensive as the number of candidates increases. To address this, Fast Best-of-N Sun et al. (2024a) follows a speculative rejection scheme. It queries the reward model multiple times throughout the generation of one candidate response and terminates the unpromising generation early based on the rewards. As such, this way accelerates the inference process without completely generating all candidates. Jinnai et al. (2024) investigate the reward backing problem of the generate-then-rank strategy. Brown et al. (2024) systematically investigate the scaling of testing-time computing with various ORMs across multiple tasks.
Ranking by Process Rewards.
As aforementioned, outcome reward models may struggle to discern the nuance among candidate responses. Thus many methods adopt process reward models (PRMs) for the generate-then-rank strategy. These methods score intermediate steps of candidate responses through a PRM and aggregate these step-level scores through multiplication or minimum to compute an overall score for ranking or voting Zhang et al. (2025g).
Early work by Lightman et al. (2023) introduces a PRM trained on a large-scale human-labeled math dataset (PRM800K) and ranks candidate math solutions by the product of their step-level reward scores. DIVERSE Li et al. (2023b) trains a reward model to assign scores to both the entire path and individual reasoning steps. Then it picks up the best one from multiple candidates through weighted voting. Math-Shepherd Wang et al. (2023b) uses a PRM to score each step in a math solution and ranks solutions according to the lowest step-level score. Notably these methods can improve reasoning consistency—the chosen solution builds on a series of reliable steps rather than merely delivering a correct final answer.
Some studies have explored multimodal PRMs for the generate-then-rank strategy. Examples include VisualPRM Wang et al. (2025c) and ViLPRM Tu et al. (2025). They incorporate both image and text inputs with step-wise evaluation in the multimodal reasoning. Some approaches combine outcome and process rewards to improve ranking quality. For example, CoRe Zhu et al. (2022) integrates these two signals to jointly verify model outputs during inference.
4.2 Reward-Guided Decoding
While the above generate-then-rank approach is simple and effective, it inherently decouples generation from evaluation, limiting its ability to refine outputs dynamically during decoding. In contrast, reward-guided decoding tightly incorporates reward signals to guide the generation process of language models. Based on the granularity of guidance, we categorize this line of work into two strategies: token-level guidance and step-level guidance. As shown in Figure 8(c), these strategies guide the language model’s token-level or step-level decoding based on the reward signals through a search algorithm, such as greedy search, beam search, or Monte Carlo Tree Search (MCTS). This enables fine-grained control over output quality and alignment and can foster reasoning and planning abilities.
Token-level Guidance.
Token-level guidance steers language model generation by incorporating reward signals into the token selection process at each decoding step. This strategy commonly combines the token likelihoods with the reward signals from a reward model.
RAD Deng and Raffel (2023) adjusts token selection by combining the token’s likelihood and the scalar rewards. It can control output attributes such as non-toxicity and sentiment. ARGS Khanov et al. (2024) applies a similar strategy to align the helpfulness and harmless preferences. This work converts preference alignment from the previous training stage to the inference stage. PG-TD Zhang et al. (2023c) targets code generation. It uses the pass rate over test cases as reward signals and integrates them into MCTS to guide token-level planning and generation and integrates it into MCTS to guide token-level planning and generation.
However, the above methods are limited by low decoding efficiency since they require querying the reward model for each candidate token at every decoding step. To improve decoding efficiency, ARM Troshin et al. (2024) proposes a low-rank approximation of the reward model to score all token candidates within a single query. Similarly, FaRMA Rashid et al. (2025) trains a reward model that scores all token candidates in a single forward pass.
Step-level Guidance.
Beyond token-level guidance, step-level guidance operates on intermediate steps of the generation process. As illustrated in Figure 6(d), the generation is decomposed into multiple intermediate steps. At each step, a search algorithm, such as beam search and Monte Carlo Tree Search (MCTS), explores the output space and selects the appropriate steps with the guidance of reward signals. This mechanism enables the model to recover from earlier errors and enhance reasoning.
In some cases, a step may correspond to a semantic segment. For example, CARDS Li et al. (2024a) samples candidate semantic segments at each step. Then it uses a reward model to score the resulting output with a segment and choose the high-reward segment for continuing decoding.
More studies focus on guiding reasoning steps with reward signals. GRACE Khalifa et al. (2023) trains a discriminator as the reward model that scores the correctness of candidate reasoning steps. During decoding, it combines the step-level reward scores with the language model’s likelihood to guide step selection toward more accurate reasoning paths. Xie et al. (2023) employs a similar approach by prompting the language model itself to evaluate the reasoning steps during beam search. Snell et al. (2025) use a process reward model to guide which reasoning steps are retained in the beam search and look-ahead search during decoding. ORPS Yu et al. (2024c) derives outcome rewards from code execution feedback and process rewards from LLM-generated self-critiques about code reasoning quality. It combines them to guide a tree-based search process, e.g., selecting, expanding, and refining candidate solutions throughout the generation. RSD Liao et al. (2025) combines rewards with speculative decoding. During its speculative decoding, it leverages a process reward model to determine whether to accept a draft model’s output or invoke the target model.
Some studies guide the decoding based on the step-level value, i.e., cumulative future rewards. Tree-of-Thoughts Yao et al. (2023) prompts the language model to assess the value of the current state by producing a scalar score or short phrases. It then guides the search algorithms, like BFS and DFS, to explore diverse reasoning trajectories. OVM Yu et al. (2023a) trains a value model to estimate the probability that a partial reasoning path leads to a correct final answer. During inference, it guides the decoding via value-based beam search to select the most promising reasoning trajectories.
Other methods use reward signals to guide Monte Carlo Tree Search (MCTS). RAP Hao et al. (2023) defines task-specific reward functions for MCTS. This encourages the model to simulate future states and select high-reward reasoning paths for planning, math, and logic tasks. STILL-1 Jiang et al. (2024) follows a similar reward-guided MCTS way. rStar Qi et al. (2024) assigns rewards to reasoning trajectories by combining the correctness of the final answer with self-consistency confidence. Then it back-propagates these rewards to the steps in the trajectories to guide future MCTS exploration toward more promising reasoning paths. Several extensions leverage process reward models to precisely guide MCTS, including ReST-MCTS* Zhang et al. (2024a), LE-MCTS Park et al. (2024), and rStar-Math Guan et al. (2025). They use step-level scoring from process reward models to select and expand high-quality reasoning trajectories during the search.
Figure 8: Illustration of Post-Inference with Rewards. (a): Self-Correction, using the language model itself. (b): Correction with External Feedback, such as trained model, external knowledge, and external tools.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=black, rounded corners, align=center, text centered, minimum width=4em, edge+=darkgray, line width=0.5pt, s sep=8pt, line width=0.5pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, minimum width=15em, fill=gray!10, leaf/.style=font=, align=left, draw=none, inner xsep=8pt, leaf2/.style=text width=10em, font=, draw=none, inner xsep=8pt , where level=1text width=6em,align=center,font=, where level=2text width=8.5em, align=center, font=, where level=3text width=17.5em, align=left,font=, where level=4align=left,font=, [ Post-Inference with Rewards, ver [ Self-Correction
§ 5.1, fill=post-inference-fill, draw=post-inference-border [ Self-Refine Madaan et al. (2023); Reflexion Shinn et al. (2023);
CoVe Dhuliawala et al. (2023); SCoRE Kumar et al. (2024);
RISE Qu et al. (2024) ,leaf, fill=post-inference-fill, draw=none, text width=17.5em ] ] [ Correction with
External Feedback
§ 5.2, fill=post-inference-fill, draw=post-inference-border [ Trained Models, fill=post-inference-fill, draw=post-inference-border [ CodeRL Le et al. (2022); RL4F Akyurek et al. (2023);
Shepherd Wang et al. (2023c); A2R Lee et al. (2024);
CTRL Xie et al. (2025b); CriticGPT McAleese et al. (2024);
DARS Li et al. (2025a); REFINER Paul et al. (2023);
AutoMathCritique Xi et al. (2024); MAD Liang et al. (2023);
Cohen et al. (2023); Du et al. (2023) ,leaf, fill=post-inference-fill, draw=none ] ] [ External Knowledge, fill=post-inference-fill, draw=post-inference-border [ RARR Gao et al. (2022); ReFeed Yu et al. (2023c);
LLM-Augmenter Peng et al. (2023); Varshney et al. (2023);
FACTOOL Chern et al. (2023) ,leaf, fill=post-inference-fill, draw=none ] ] [ External Tools, fill=post-inference-fill, draw=post-inference-border [ Self-Edit Zhang et al. (2023a); Self-Debug Chen et al. (2023);
CYCLE Ding et al. (2024); Logic-LM Pan et al. (2023a);
IHR Qiu et al. (2023); Baldur First et al. (2023);
CRITIC Gou et al. (2023); RCI Kim et al. (2023) ,leaf, fill=post-inference-fill, draw=none ] ] ] ]
Figure 9: Overview of Post-Inference with Rewards.
5 Post-Inference with Rewards
Post-inference with rewards aims to correct and refine the model outputs after they have been generated with reward signals. This approach enables iterative enhancement without updating model parameters, offering a lightweight and compatible mechanism for test-time scaling. Compared to sparse scalar rewards, post-inference methods favor critique rewards that provide fine-grained and dense signals for correction. These critiques typically identify error locations, explain reasoning flaws, and suggest revisions. To leverage such rich feedback, the language model commonly incorporates them as augmented context and iteratively generates revised outputs. According to the source of rewards, we categorize these methods into two main strategies: Self-Correction and Correction with external rewards.
5.1 Self-Correction
As illustrated in Figure 8(a), self-correction leverages the language model itself as a generative reward model to evaluate and revise its own outputs, similar to the aforementioned self-rewarding strategy. Self-Refine Madaan et al. (2023) prompts the language model itself to produce natural language feedback on its own outputs. It then leverages this feedback as reward signals to refine outputs iteratively. Similarly, Reflexion Shinn et al. (2023) generates reflection feedback through the language model itself. It additionally maintains a memory bank to store previous feedback, outputs, and scalar feedback from evaluation metrics. These reflections serve as auxiliary contexts to refine subsequent generations. CoVe Dhuliawala et al. (2023) prompts the language model to generate and answer verification questions about its own outputs to identify factual errors. This verification feedback guides the model to correct its responses to reduce hallucination.
In addition, SCoRE Kumar et al. (2024) trains the language model via reinforcement learning to enhance its self-correction capability. Similarly, RISE Qu et al. (2024) bootstraps the language model with DPO training for better self-correction.
5.2 Correction with External Feedback
While self-correction is simple, prior studies argue that general language models struggle to identify and correct their errors without external feedback, especially for less capable small-scale models Huang et al. (2023); Kamoi et al. (2024); Madaan et al. (2023); Pan et al. (2023b). Owing to this, increasing attention has been devoted to incorporating external feedback as reward signals to refine model outputs as shown in Figure 8(b). We classify these works according to the feedback source: Trained Models, External Knowledge, and External Tools.
Trained Model.
Many methods rely on more capable trained models (commonly referred to as critic models) to provide feedback as reward signals. The feedback is mostly natural language critiques containing quality assessments and correction suggestions on model outputs. The early work CodeRL Le et al. (2022) uses a trained critic model to predict the functional correctness of the generated code. Afterward, this feedback serves as reward signals to guide the language model to regenerate the code. Welleck et al. (2022) trains a critic model to provide both scalar and critique feedback for mathematics, constrained generation, and toxicity control. RL4F Akyurek et al. (2023) uses a critic model to generate critiques for topic-based summarization, action planning, and alphabetization tasks. Shepherd Wang et al. (2023c) presents a critic model that can identify factual, logical, or stylistic issues and suggest actionable refinements for the language model. A2R Lee et al. (2024) incorporates factual critiques from a critic model to mitigate the hallucination issue. CTRL Xie et al. (2025b) focuses on code generation. It fine-tunes a critic model with code execution results via GRPO, and the model provides actionable critiques to guide the refinement of generated code. McAleese et al. (2024) use a trained critic model, CriticGPT, to identify the flaws in generated code. It can output structured critiques that disclose bugs and reasoning errors in the generated code. DARS Li et al. (2025a) refines model outputs by guiding iterative reasoning with the reflection critiques from a trained critic model.
Moreover, some studies focus on fine-grained step-level feedback for correction. REFINER Paul et al. (2023) adopts a critic model to provide fine-grained feedback on the intermediate reasoning steps. It then uses these feedback to iteratively refine the language model’s reasoning for math and moral story. AutoMathCritique Xi et al. (2024) similarly trains a critic model to provide process-level critiques, which supports iterative refinement of mathematical reasoning outputs.
Some methods follow the multi-agent debate design, where critiques from peer agents support reflection and improvement. These methods include MAD Liang et al. (2023), Cohen et al. (2023), and Du et al. (2023).
In the multimodal field, DRESS Chen et al. (2024b) leverages GPT-4 to generate feedback on vision-language model outputs, including critiques and refinement suggestions. It then uses these feedback to guide the model to refine its outputs iteratively.
External Knowledge.
External knowledge sources mainly provide factual critiques along with retrieved evidence, which can improve factuality and reduce hallucinations. RARR Gao et al. (2022) derives hybrid rewards based on the entailment-based agreement between the model output and retrieved evidence from external knowledge. These reward signals then guide the post-hoc correction of the language model to improve factual attribution while preserving the original text’s intent and structure. ReFeed Yu et al. (2023c) applies a similar method to knowledge-intensive QA tasks. LLM-Augmenter Peng et al. (2023) computes the Knowledge F1 scores between model outputs and retrieved evidence from external knowledge as reward signals. These signals guide the language model’s decision to either finalize or continue to refine its response. Varshney et al. (2023) formulates verification questions regarding the low-confidence concepts in the model outputs. To verify these questions, it retrieves evidence from external knowledge and takes the results and evidence as feedback to guide the language model’s refinement and reduce hallucination. As a general factuality tool, FACTOOL Chern et al. (2023) broadens this idea to an enormous scope, including knowledge-based QA, code generation, mathematical reasoning, and scientific literature review.
External Tools.
External tools can execute and verify the language model outputs, and their feedback can work as reward signals for correction. Self-Edit Zhang et al. (2023a) and Self-Evolve Jiang et al. (2023) use program execution feedback from code compilers to guide the refinement of the language model. Self-Debug Chen et al. (2023) and CYCLE Ding et al. (2024) extend them with more feedback, for instance, unit test results and program explanations.
Apart from code compilers, Logic-LM Pan et al. (2023a) uses the feedback from symbolic logic reasoner as critique rewards to refine the model’s answers to logic reasoning problems. IHR Qiu et al. (2023) depends on the feedback from a symbolic interpreter for inductive reasoning, and Baldur First et al. (2023) incorporates feedback from a proof checker for automated formal verification. CRITIC Gou et al. (2023) and RCI Kim et al. (2023) can leverage feedback from diverse external tools, such as search engines, code interpreters, and calculators.
Figure 10: Illustration of Benchmarking Reward Models. Annotations come from human annotators or AI annotators with human verification. The annotations are mainly pointwise (e.g., a scalar score for each sample) or pairwise (e.g., chosen and rejected responses).
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=black, rounded corners, align=center, text centered, minimum width=4em, edge+=darkgray, line width=0.5pt, s sep=8pt, line width=0.5pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, minimum width=18em, fill=gray!10, leaf/.style=font=, align=left, draw=none, inner xsep=8pt, leaf2/.style=text width=10em, font=, draw=none, inner xsep=8pt , where level=1text width=10em,align=center,font=, where level=2text width=21.5em, align=center, font=, where level=3text width=17.5em, align=left,font=, where level=4align=left,font=, [ Benchmarking Reward Models, ver [ Benchmarking
Outcome Reward Models
§ 6.1, fill=benchmark-fill, draw=benchmark-border [ Zheng et al. (2023); RewardBench Lambert et al. (2024);
RM-Bench Liu et al. (2024d); AceMath-RewardBench Liu et al. (2024e);
RMB Zhou et al. (2024a); CriticBench Lin et al. (2024b);
MetaCritique Sun et al. (2024b) ,leaf, fill=benchmark-fill, draw=none ] ] [ Benchmarking
Process Reward Models
§ 6.2, fill=benchmark-fill, draw=benchmark-border [ MathCheck-GSM Zhou et al. (2024d); MR-GSM8K Zeng et al. (2023);
ProcessBench Zheng et al. (2024); PRMBench Song et al. (2025b);
Big-Bench Mistake Tyen et al. (2023); MR-Ben Zeng et al. (2024) ,leaf, fill=benchmark-fill, draw=none ] ] [ Benchmarking
Multimodal Reward Models
§ 6.3, fill=benchmark-fill, draw=benchmark-border [ MJ-Bench Chen et al. (2024c); MLLM-as-a-Judge Chen et al. (2024a);
VL-RewardBench Li et al. (2024b);
Multimodal-RewardBench Yasunaga et al. (2025) ;
SVIP Gao et al. (2025); VLRMBench Ruan et al. (2025) ,leaf, fill=benchmark-fill, draw=none ] ] [ Other Benchmarks
§ 6.4, fill=benchmark-fill, draw=benchmark-border [ RAG-RewardBench Jin et al. (2024); M-RewardBench Gureja et al. (2024);
PPE Frick et al. (2024) ,leaf, fill=benchmark-fill, draw=none ] ] ]
Figure 11: Overviews of Benchmarking Reward Models.
6 Benchmarking Reward Models
As reward models play a central role in the learning-from-rewards paradigm for post-training and test-time scaling, rigorous and diverse benchmarks are essential for evaluating their capabilities, like other benchmarks for LLMs Pan et al. (2024b, 2023c); Zhou et al. (2024b); Wu et al. (2022, 2023a, 2023b, 2024c, 2024d, 2024e, 2024b, 2024f, 2024g); Wu (2024). As illustrated in Figure 10, recent benchmarking efforts primarily rely on expert human annotators or AI annotators (e.g., LLM-as-a-Judge frameworks) followed by human verification to ensure reliability. The resulting annotations are mainly pointwise (e.g., scalar scoring) or pairwise (e.g., selecting the preferred response given two options). These benchmarks vary in task coverage, evaluation protocols, annotation sources, and reward formats. We discuss the representative benchmarks in the literature as follows.
6.1 Benchmarking Outcome Reward Models
A dominant line of benchmarking studies centers on outcome reward models that evaluate the overall quality of generated outputs. Zheng et al. (2023) is an early work that evaluates LLMs’ judging ability by directly prompting them. As LLMs can naturally function as generative reward models, this study also represents one of the earliest benchmarks for reward models. RewardBench Lambert et al. (2024) is the first comprehensive benchmarks for reward models. It aggregates preference data from existing datasets, such as AlpacaEval and MTBench, to evaluate reward model performance in chatting, reasoning, and safety. RM-Bench Liu et al. (2024d) introduces evaluation for reward models on sensitivity to subtle content changes and robustness to style biases. It constructs preference pairs across chat, code, math, and safety domains using GPT-4o. AceMath-RewardBench Liu et al. (2024e) focuses on math-specific evaluations. It tests whether reward models can identify correct solutions from candidates across various mathematical tasks and difficulty levels. RMB Zhou et al. (2024a) furthermore broadens the evaluation scope to 49 real-world scenarios.
Apart from evaluating with preference data, some benchmarks focus on the critique ability of reward models. CriticBench Lin et al. (2024b) assess whether reward models can generate critiques that accurately identify the correctness of a response and effectively guide the correction. Similarly, MetaCritique Sun et al. (2024b) benchmarks LLM-generated critiques by decomposing them into atomic information units and assessing their correctness.
6.2 Benchmarking Process Reward Models
Recently more benchmarks focus on process reward models due to their increasing significance. In detail, several benchmarks focus on math reasoning, such as MathCheck-GSM Zhou et al. (2024d), MR-GSM8K Zeng et al. (2023), and MR-MATH Xia et al. (2024). They require reward models to locate the first error step in a math reasoning solution. Their testing samples are adapted from existing math datasets, including GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021). Furthermore, ProcessBench Zheng et al. (2024) features diversity and higher difficulty levels by scaling this up to Olympiad- and competition-level math problems He et al. (2024); Gao et al. (2024b). Beyond step correctness, PRMBench Song et al. (2025b) offers a more fine-grained benchmark. It annotates each step in the reasoning path with specific error types grouped into three dimensions: simplicity, soundness, and sensitivity. The annotations come from LLM-generated perturbations and are subsequently verified by human annotators.
Besides mathematical reasoning, Big-Bench Mistake Tyen et al. (2023) targets logical reasoning. It annotates chain-of-thought trajectories from BIG-Bench bench authors (2023), each labeled with the first logical error. Furthermore, MR-Ben Zeng et al. (2024) expands this to the reasoning process of seven domains: math, logic, physics, chemistry, medicine, biology and code.
6.3 Benchmarking Multimodal Reward Models
Due to the prevalence of multimodal language models, another vital line of benchmarks focuses on multimodal reward models with diverse evaluation protocols.
MJ-Bench Chen et al. (2024c) depends on text-to-image generation tasks for evaluation. It builds preference data across four dimensions: text-image alignment, safety, image quality, and social bias. MLLM-as-a-Judge Chen et al. (2024a) uses image understanding tasks for benchmarking and includes pointwise and pairwise scoring. VL-RewardBench Li et al. (2024b) includes three tasks: general multimodal instructions, hallucination detection, and multimodal reasoning. Multimodal-RewardBench Yasunaga et al. (2025) spans six key capabilities of multimodal reward models: general correctness, human preference, factual knowledge, reasoning, safety, and VQA.
Beyond the outcome level, current benchmarks also assess multimodal process reward models. SVIP Gao et al. (2025) targets process-level evaluation on relevance, logic, and attribute correctness of diverse multimodal tasks. It transforms reasoning paths into executable visual programs and automatically annotates each step. VLRMBench Ruan et al. (2025) further integrates evaluation on three dimensions: reasoning steps, whole outcomes, and critiques on error analysis. It collects testing data of multimodal understanding through AI annotations and human verification.
6.4 Other Benchmarks
In addition to general-purpose evaluations, several benchmarks aim to address domain-specific or emerging challenges in reward modeling. RAG-RewardBench Jin et al. (2024) targets reward model evaluation in RAG. It constructs preference data for RAG-specific scenarios, including multi-hop reasoning, fine-grained citation, appropriate abstention, and conflict robustness. M-RewardBench Gureja et al. (2024) extends the evaluation to multilingual contexts. Instead of direct evaluation, PPE Frick et al. (2024) indirectly evaluates reward models through RLHF pipelines. It measures the performance of trained LLMs with a reward model, offering a practical perspective.
7 Applications
The strategies described above for learning from rewards have been widely adopted across diverse applications. Early applications focus on preference alignment, such as RLHF Ouyang et al. (2022) and RLAIF Bai et al. (2022b). In particular, the recent DeepSeek-R1 DeepSeek-AI et al. (2025) has demonstrated the effectiveness of reinforcement learning to develop large reasoning models, which has inspired a wave of R-1 style applications for diverse areas. In this section, we review the primary applications following these strategies.
7.1 Preference Alignment
Learning-from-rewards strategies have become the cornerstone for aligning LLMs with human preferences. These strategies design diverse reward signals to encourage desirable attributes, such as factuality, harmlessness, and helpfulness, while penalizing undesired behaviors like toxicity, bias, and hallucination. We summarize three major objectives of preference alignment as follows.
  • Factuality and Reducing hallucination. Hallucination refers to generating fluent but factually incorrect or fabricated content Tian et al. (2023). It is a pervasive issue for language models, especially in knowledge-intensive tasks such as healthcare and scientific research. The methods for this alignment span the training, inference, and post-inference stages Sun et al. (2023); Lin et al. (2024a); Zhao et al. (2023); Peng et al. (2023); Wang et al. (2023c). The rewards mainly stem from human preferences about factuality as well as external knowledge sources. For instance, Fact-RLHF Sun et al. (2023) trains a factuality-aware reward model on human preferences and additional supervision from image captions and multiple-choice answers The reward model is then used to fine-tune the multimodal language model via PPO, guiding the model to reduce hallucinations. RLFH Wen et al. (2024) decomposes the model responses into atomic statements, verifies their truthfulness against external knowledge, and converts them into dense token-level scalar rewards. To reduce hallucination, it directly uses these reward signals to fine-tune the model via PPO.
  • Safety and Harmlessness. Safety and harmlessness constitute another critical axis of alignment, particularly in adversarial or socially sensitive contexts Bai et al. (2022b); Ji et al. (2023). Language models must be discouraged from producing toxic, offensive, or biased content before being deployed in real-world systems. To this end, the methods primarily focus on the training Ouyang et al. (2022); Bai et al. (2022a) and inference stages Deng and Raffel (2023); Khanov et al. (2024). For instance, RAD Deng and Raffel (2023) depends on reward signals to produce non-toxicity content during decoding.
  • Helpfulness. Meanwhile, helpfulness emphasizes that language models are expected to provide relevant, informative, and context-aware responses to fulfill user intent Taori et al. (2023). This alignment is imperative in areas like instruction-following and dialogue systems. Reward signals are generally sourced from human preferences and task-specific quality metrics Bai et al. (2022a).
7.2 Mathematical Reasoning
Mathematical reasoning is vital to measure the language model’s ability to solve complex reasoning problems. Some methods build reward models and fine-tune the language model for math reasoning Shao et al. (2024); DeepSeek-AI (2025), particularly using process reward models Uesato et al. (2022); Luo et al. (2023) like Math-Shepherd Wang et al. (2023b). They can provide step-level reward signals for a math reasoning solution. Moreover, some approaches construct preference data for math reasoning, i.e., correct and incorrect solutions, and then fine-tune the language model through DPO Lai et al. (2024); Xu et al. (2025a). Others include inference-time scaling strategies, such as generate-then-rank Cobbe et al. (2021); Lightman et al. (2023), and reward-guided decoding with search algorithms like MCTS Hao et al. (2023); Guan et al. (2025).
7.3 Code Generation
The code generation task has made significant strides due to the development of LLMs, which improves software engineering productivity by a large margin. To improve the code language model through fine-tuning, the reward signals can come from various sources, including Zhu et al. (2024), and code compiler feedback, unit test results, and code analysis Liu et al. (2023); Dou et al. (2024); Tao et al. (2024); Zhou et al. (2025a). For example, DeepSeek-Coder-V2 Zhu et al. (2024) trains a reward model for code generation and fine-tunes the language model via the reinforcement learning algorithm GRPO Shao et al. (2024). Additionally, some approaches guide the inference of language models during code generation with reward models, including the generate-then-rank Ni et al. (2023); Hosseini et al. (2024) and reward-guided decoding Yu et al. (2024c). Another popular direction refines the generated code to correct errors and bugs through the language model itself Shinn et al. (2023); Zhang et al. (2023a); Chen et al. (2023) or external feedback Xie et al. (2025b).
7.4 Multimodal Tasks
Learning-from-rewards strategies have been widely applied to multimodal tasks, including multimodal understanding and generation. Most studies adopt reinforcement learning and reward-guided decoding methods. For instance, Q-Insight Li et al. (2025d) focuses on improving comprehensive image quality understanding with reinforcement learning. VLM-R1 Shen et al. (2025a) applies reinforcement learning to fine-tune vision-language models and focuses on two tasks: referring expression compression and object detection. Vision-R1 Huang et al. (2025) enhances multimodal reasoning of vision-language models for mathematical VQA. Zhan et al. (2025) proposes another Vision-R1, but it mainly facilitates object localization tasks with vision-language models.
Video-R1 Feng et al. (2025b), VideoChat-R1 Li et al. (2025f), and TinyLLaVA-Video-R1 Zhang et al. (2025e) apply GRPO into video reasoning. R1-V Chen et al. (2025a) and CrowdVLM-R1 Wang et al. (2025e) focus on visual counting. More example applications include multimodal reasoning Zhou et al. (2025b); Meng et al. (2025); Tan et al. (2025); Li et al. (2025b); Liu et al. (2025f), object detection Liu et al. (2025h), segmentation Liu et al. (2025d), and image/video generation Guo et al. (2025b); Liu et al. (2025a).
7.5 Agents
LLM Agent is an autonomous system that automatically performs complex tasks through task decomposition and action execution in dynamic environments Wang et al. (2024c). Various learning-from-rewards strategies have been applied to training or guiding the agents. AgentRM Xia et al. (2025) targets general-purpose decision-making agents across domains such as web navigation, embodied planning, text games, and tool use. During inference, a reward model guides the agents to choose candidate actions or trajectories. AgentPRM Choudhury (2025) trains LLM agents with a process reward model. KBQA-o1 Luo et al. (2025) guides MCTS with a reward model for the knowledge base question answering task with agents. DeepResearch OpenAI (2025) and DeepResearcher Zheng et al. (2025) design agents for research tasks. They both use reinforcement learning to fine-tune the agents. UI-R1 Lu et al. (2025) introduces a rule-based reinforcement learning framework for GUI action prediction with multimodal agents. InfiGUI-R1 Liu et al. (2025c) is a similar work with GUI agents. RAGEN Wang et al. (2025f) propose training the agents via multi-turn reinforcement learning with a new algorithm based on GRPO.
7.6 Other Applications
Many other applications have been developed following the learning-from-rewards strategies.
Embodied AI is essential for the development of artificial general intelligence. AI systems, such as embodied robots, must interact with the physical world and complete complex tasks through high-level planning and low-level control. They generally aim to enhance the embodied reasoning abilities with reinforcement learning, such as Cosmos-reason1 Azzolini et al. (2025), iRe-VLA Guo et al. (2025a), Embodied-Reasoner Zhang et al. (2025d), and Embodied-R Zhao et al. (2025a).
Several approaches apply reinforcement learning to reason with information retrieval from knowledge databases or the real-world web. These approaches include R1-Searcher Song et al. (2025a), Search-R1 Jin et al. (2025), DeepRetrieval Jiang et al. (2025), ReSearch Chen et al. (2025b), and WebThinker Li et al. (2025e). They adopt different reward designs to improve search performance.
Applications for other applications also emerge. ToRL Li et al. (2025g), ReTool Feng et al. (2025a), SWi-RL Goldie et al. (2025), ToolRL Qian et al. (2025) and OTC Wang et al. (2025a) are proposed to improve LLMs’ reasoning ability to call various tools through reinforcement learning. Rec-R1 Lin et al. (2025) applies reinforcement learning for recommendation system. SWE-RL Wei et al. (2025) aims at software engineering with reinforcement learning. SQL-R1 Ma et al. (2025) focuses on natural language to SQL reasoning. It uses a composite reward function with format correctness, execution success, result accuracy, and reasoning completeness.
Some applications are designed for specific areas. Med-R1 Lai et al. (2025) and MedVLM-R1 Pan et al. (2025) are proposed for medical field. They target medical VQA across various imaging modalities (e.g., CT, MRT, and X-ray) and several clinical tasks, like diagnosis, and anatomy identification. Fin-R1 Liu et al. (2025e) develops LLMs for the financial field, targeting financial QA and decision-making. It leverages accuracy and format rule-based rewards to train a language model on domain-specific data. DianJin-R1 Zhu et al. (2025) is another LLM for the financial field with reinforcement learning.
Figure 12: Illustration of challenges and future directions.
8 Challenges and Future Directions
In this section, we discuss the current challenges and future directions of learning from rewards. Figure 12 summarizes the key challenges and future directions from the perspective of reward model design and learning strategies. Ultimately, we envision the development of interpretable, robust, and continually evolving agent systems capable of interacting with and adapting to the complexities of the real world.
8.1 Interpretability of Reward Models
Interpretability of reward models remains an open challenge for the learning-from-rewards strategies Russell and Santos (2019); Zhang et al. (2023d); Jenner and Gleave (2022). Most reward models are typically treated as black boxes that produce scalars or critiques without exposing human-interpretable explanations. Such opacity hinders human trust and oversight and may lead to misaligned optimization. In consequence, enhancing reward model interpretability is essential for reliable alignment, enabling humans to inspect and verify the internal decision process and steer models toward desired behavior. Recent efforts have attempted to address this issue. For instance, ArmoRM Wang et al. (2024b) improves the interpretability with multi-objective reward modeling, where each objective corresponds to a human-interpretable dimension, such as helpfulness, correctness, coherence, complexity, and verbosity. While this approach is effective, its interpretability is limited to these predefined objectives. In addition, emerging generative reward models can disclose their rationales of reward scoring Zhao et al. (2025b); Khalifa et al. (2025). While promising, their interpretability remains limited and demands further investigation into consistency, reliability, and faithfulness.
8.2 Generalist Reward Models
A promising future direction is the development of generalist reward models. Most existing reward models are designed for narrow domains; thus they often suffer from weak generalization across tasks. Moreover, their reward outputs are typically static and lack support for inference-time scalability, hindering their application in diverse and open-ended scenarios Liu et al. (2024a); Zhang et al. (2024c); Snell et al. (2025).
In contrast, a generalist reward model seeks to overcome these limitations. They demand flexibility for input types, including single, paired, or multiple responses, and also require accurate reward generation in various domains, such as question answering, math reasoning, and code generation. Besides, they are expected to generate higher-quality reward signals with increased inference-time computing. Such models offer a unified interface for reward modeling across domains and enable scalable, interpretable reward generation. For example, DeepSeek-GRM Liu et al. (2025g), a recent attempt in this direction, proposes a pointwise generative reward model. Rather than only scalars, it can generate evaluative natural language principles and critiques, enabling effective inference-time scaling through multi-sample voting and meta-reward filtering.
8.3 Reward Hacking
Reward hacking is a fundamental challenge in learning from rewards Everitt et al. (2021); Amodei et al. (2016); Weng (2024); Liu et al. (2025b). It occurs when models exploit unintended shortcuts in the reward function to obtain high rewards without truly learning the desired behaviors or completing the task as designed. This phenomenon has been observed across domains. For instance, LLMs may fabricate plausible yet incorrect answers, and code LLMs subtly modify unit tests to pass evaluations Denison et al. (2024). Reward backing can also happen during inference, called in-context reward hacking Pan et al. (2024c, a). It arises in self-refinement loops where the same model acts as both the generator and the judge. In such cases, the model may learn to produce outputs that exploit its own evaluation heuristics, leading to inflated internal scores while deviating from true objectives.
Reward hacking fundamentally arises from the difficulty of specifying a reward function that perfectly captures the true objectives. As articulated by Goodhart’s Laws—When a measure becomes a target, it ceases to be a good measure—any proxy metric used as a reward will eventually be exploited once applying optimization pressure. To mitigate reward hacking, the following directions are worth exploring: (i) Designing more robust and tamper-resistant reward functions Razin et al. (2025); Shen et al. (2025b); Peng et al. (2025); (ii) Detecting misalignment via behavioral or distributional anomaly detection Pan et al. (2022); (iii) Decoupling feedback mechanisms to prevent contamination Uesato et al. (2020); (iv) Auditing the dataset for training reward models to reduce reward hacking risks Revel et al. (2025).
8.4 Grounded Rewards from Real-World Interactions
Despite recent advances in learning from rewards for LLMs, most methods fundamentally rely on human preferences or well-curated automated feedback. The LLMs are typically optimized to maximize the rewards derived from these feedback. In consequence, this inherently limits the agent’s ability to surpass existing human knowledge and adapt to complex environments.
Due to these limitations, moving beyond chat-driven rewards toward grounded real-world rewards is another promising direction. This movement requires LLMs to be integrated into agentic frameworks, and agents should increasingly interact directly with their environment and derive reward signals from observed outcomes. For example, a health assistant could optimize behavior based on physiological signals rather than user ratings, and a scientific agent could refine hypotheses based on experimental data rather than expert approval Silver and Sutton (2025). This shift would enable agents to close the feedback loop with the real world, allowing for autonomous discovery, adaptation, and pursuit of goals beyond human understanding. The transition to real-world interactions raises substantial technical challenges. Agents must handle noisy, delayed, or partial feedback from complex environments, requiring advances in credit assignment, robust exploration, and uncertainty modeling.
8.5 Continual Learning from Rewards
Current learning-from-rewards strategies often assume a fixed dataset, a predefined reward model, and short episodic interactions. Once trained, models typically exhibit limited abilities to adapt to new tasks or evolving environments Zhang et al. (2024b); Silver and Sutton (2025). This episodic and offline paradigm sharply contrasts with real-world intelligence’s dynamic, ongoing nature, where agents must continually learn from experience and recalibrate based on new feedback.
As such, a vital direction is continual learning from rewards. It is a crucial foundation for building lifelong competent and aligned agents. By abandoning the traditional assumption of fixed objectives, models can remain responsive to changing reward signals, avoid performance degradation under distributional shifts, and better reflect long-term user intent. Notably, it is a broader idea of continual reinforcement learning Abel et al. (2023); Li et al. (2024c); Bowling and Elelimy (2025). Achieving continual learning from rewards presents significant challenges. It requires addressing catastrophic forgetting, maintaining stability while enabling plasticity, and designing dynamic reward modeling mechanisms.
9 Conclusion
In this paper, we present a comprehensive survey of learning LLMs from rewards. We categorize the landscape into three key stages-training, inference, and post-inference, each reflecting a distinct paradigm for integrating reward signals into steering LLMs’ behavior. For each stage, we review representative strategies in terms of reward signal design and learning algorithms. In addition, we summarize recent progress in benchmarking reward models and applications. Finally we identify core challenges and outline promising future directions. We hope this survey provides a structured understanding of the field and inspires future research.